What is the gravitational field strength at the surface of a planet if a 2, k, g,


2kg basket of shopping weighs 52, N,


52N?


1
Expert's answer
2021-11-19T10:49:14-0500

We can find the gravitational field strength at the surface of the planet from the definition of the weight:


"W=mg,""g=\dfrac{W}{m}=\dfrac{52\ N}{2\ kg}=26\ \dfrac{N}{kg}."